# Revisiting the Equivalence of Bayesian Neural Networks and Gaussian Processes: On the Importance of Learning Activations

### v2

@@ B.5 Periodic and Conditional Activations
Appendix C Extensive comparison against Tran et al., (2022)
@@
A related approach, formulated as the dual problem of optimal transport, was explored in (Tran et al., , 2022). However, their method utilizes the complex Normalizing Flow priors as default, and as such, does not incorporate any mathematical property that guarantees the resulting model is a GP. From this perspective, the approach in (Tran et al., , 2022) can be seen as a heuristic that yields favorable performance -- measured by RMSE and NLL -- within the training data range, but without any formal assurance that the model ultimately corresponds to a GP. Their method relies on a multiple-layer BNN with narrow layers, where the number of layers is chosen empirically based on the specific problem. Although employing a Gaussian prior may encourage GP-like behavior, this property does not necessarily hold when using Normalizing Flow priors.
@@
C.1 Usage of wider BNNs
First, we investigated the effect of increasing the width of a single-hidden-layer BNN from 50 neurons -- following the typical choice in (Tran et al., , 2022) -- to 1024 neurons. Specifically, we compared two types of priors: Gaussian priors and Normalizing Flow priors. Additionally, we examined four different activation functions commonly used in BNNs and evaluated these settings on a 1D regression task.
@@ C.1 Usage of wider BNNs
In contrast, using wider layers with Normalizing Flow priors significantly degrades performance across all considered metrics (see Tab.8). We hypothesize that this is due to the lack of theoretical foundations for such priors. Furthermore, we visualize the posterior distributions of narrow and wide BNN layers for two activation functions: ReLU, which performs poorly even with a Gaussian prior, and TanH (see Fig.9). These results indicate that Normalizing Flow priors -- despite being uniquely assigned to each neuron -- fail to capture the increased capacity of wider networks, regardless of the activation function used. This serves as a clear example of where the heuristic proposed by (Tran et al., , 2022) collapses.
Table 7: Comparison of Tran et al., (2022) using a narrow and a wide hidden layer in a BNN with a set Gaussian prior. We observe that the wider hidden layer usually helps in achieving better Wasserstein metrics and similar or slightly worse RMSE and NLL when using Gaussian priors.
@@
Table 8: Comparison of Tran et al., (2022) using a narrow and a wide hidden layer in a BNN with a set Normalizing Flow prior. We notice that using wider layers with Normalizing Flow priors significantly lowers the result on each of the considered metrics.
@@
C.2 Influence of number of hidden layers in BNNs
Next, we examine whether deeper BNNs within the (Tran et al., , 2022) framework lead to improved results.
Specifically, we compare BNNs with multiple hidden layers (i.e., $\{2,3,4,5\}$) against single-hidden-layer baselines, evaluating four distinct activation functions -- ReLU, TanH, Swish, and RBF -- on RMSE and NLL, as shown in Fig. 10.
@@ C.2 Influence of number of hidden layers in BNNs
Figure 10: Comparison of different priors, activation functions and BNNs depth in Tran et al., (2022) method in terms of RMSE (left column) and NLL right column on test data. We observe that this method with Gaussian prior (top row) usually get better results with a larger number of hidden layers. Contrary, using Normalizing Flow prior get better results only for some (worse) activations. The largest increase of quality might be observed for the worst activation functions. We find that our approach (dashed black line) has better, or similar results in terms of both NLL and RMSE.
@@
C.3 Out of distribution evaluation
Finally, we evaluate the (Tran et al., , 2022) method with different priors and varying numbers of hidden layers on an out-of-distribution (OOD) task. Since the true data distribution in OOD regions is unknown, we rely on optimal transport metrics—1-Wasserstein ($\mathcal{W}_{1}^{1}$) and 2-Wasserstein ($\mathcal{W}_{2}^{2}$). Additionally, we compare all these configurations against our single-hidden-layer method.
@@ C.3 Out of distribution evaluation
The results, presented in Fig. 13, 14, 15, and 16, lead to several key observations regarding the performance on OOD tasks:
@@
  • Our method: Due to its grounding in theoretical properties, our approach consistently outperforms all configurations of the (Tran et al., , 2022) method.
@@
Figure 12: We compare our method with Tran et al., (2022) on the out of distribution data. For this aim, we provide a wide experimental setting comparing both methods on a few different ranges of data ($x$). We present these ranges as different colour lines in this plot. Firstly, we consider the original test range (R0), then we focus on the data between the given blobs of points (R3). Finally, we consider the out of distribution ranges -- ones being closer to the train data (R2, R4, and R2 $\bigcup$ R4) or further (R1, R5, and R1 $\bigcup$ R5). The densities of evaluation points meets the density of the original data on each range despite R0.
Figure 13: Comparison of Tran et al., (2022) method with a Gaussian prior and different number of hidden layers. In this plot, we present $\mathcal{W}_{1}^{1}$ (top row) and $\mathcal{W}_{2}^{2}$ (bottom row) for the regions R3, R4 $\bigcup$ R2, and R1 $\bigcup$ R5 from left to right side. The appropriate regions might be found in Fig. 12.
Figure 14: Comparison of Tran et al., (2022) method with a Gaussian prior and different number of hidden layers. In this plot, we present $\mathcal{W}_{1}^{1}$ (top row) and $\mathcal{W}_{2}^{2}$ (bottom row) for the regions R1, R2, R4, and R5 from left to right side. The appropriate regions might be found in Fig. 12.
Figure 15: Comparison of Tran et al., (2022) method with a Normalizing Flow prior and different number of hidden layers. In this plot, we present $\mathcal{W}_{1}^{1}$ (top row) and $\mathcal{W}_{2}^{2}$ (bottom row) for the regions R3, R4 $\bigcup$ R2, and R1 $\bigcup$ R5 from left to right side. The appropriate regions might be found in Fig. 12.
Figure 16: Comparison of Tran et al., (2022) method with a Normalizing Flow prior and different number of hidden layers. In this plot, we present $\mathcal{W}_{1}^{1}$ (top row) and $\mathcal{W}_{2}^{2}$ (bottom row) for the regions R1, R2, R4, and R5 from left to right side. The appropriate regions might be found in Fig. 12.

### v3

@@ D.5 Periodic and Conditional Activations
Appendix E Extensive Comparison Against Tran et al. [2022]
@@
A related approach, formulated as the dual problem of optimal transport, was explored in [Tran et al., 2022]. However, their method utilizes the complex Normalizing Flow priors as default, and as such, does not incorporate any mathematical property that guarantees the resulting model is a GP. From this perspective, the approach in [Tran et al., 2022] can be seen as a heuristic that yields favorable performance – measured by RMSE and NLL – within the training data range, but without any formal assurance that the model ultimately corresponds to a GP. Their method relies on a multiple-layer BNN with narrow layers, where the number of layers is chosen empirically based on the specific problem. Although employing a Gaussian prior may encourage GP-like behavior, this property does not necessarily hold when using Normalizing Flow priors.
@@
E.1 Usage of Wider BNNs
First, we investigated the effect of increasing the width of a single-hidden-layer BNN from 50 neurons – following the typical choice in [Tran et al., 2022] – to 1024 neurons. Specifically, we compared two types of priors: Gaussian priors and Normalizing Flow priors. Additionally, we examined four different activation functions commonly used in BNNs and evaluated these settings on a 1D regression task.
@@ E.1 Usage of Wider BNNs
In contrast, using wider layers with Normalizing Flow priors significantly degrades performance across all considered metrics (see Tab.10). We hypothesize that this is due to the lack of theoretical foundations for such priors. Furthermore, we visualize the posterior distributions of narrow and wide BNN layers for two activation functions: ReLU, which performs poorly even with a Gaussian prior, and TanH (see Fig.13). These results indicate that Normalizing Flow priors – despite being uniquely assigned to each neuron – fail to capture the increased capacity of wider networks, regardless of the activation function used. This serves as a clear example of where the heuristic proposed by [Tran et al., 2022] collapses.
Table 9: Comparison of Tran et al. [2022] using a narrow and a wide hidden layer in a BNN with a set Gaussian prior. We observe that the wider hidden layer usually helps in achieving better Wasserstein metrics and similar or slightly worse RMSE and NLL when using Gaussian priors.
@@
Table 10: Comparison of Tran et al. [2022] using a narrow and a wide hidden layer in a BNN with a set Normalizing Flow prior. We notice that using wider layers with Normalizing Flow priors significantly lowers the result on each of the considered metrics.
@@
E.2 Influence of Number of Hidden Layers in BNNs
Next, we examine whether deeper BNNs within the [Tran et al., 2022] framework lead to improved results.
Specifically, we compare BNNs with multiple hidden layers (i.e., $\{2,3,4,5\}$) against single-hidden-layer baselines, evaluating four distinct activation functions – ReLU, TanH, Swish, and RBF – on RMSE and NLL, as shown in Fig. 14.
@@ E.2 Influence of Number of Hidden Layers in BNNs
Figure 14: Comparison of different priors, activation functions and BNNs depth in Tran et al. [2022] method in terms of RMSE (left column) and NLL right column on test data. We observe that this method with Gaussian prior (top row) usually get better results with a larger number of hidden layers. Contrary, using Normalizing Flow prior get better results only for some (worse) activations. The largest increase of quality might be observed for the worst activation functions. We find that our approach (dashed black line) has better, or similar results in terms of both NLL and RMSE.
@@
E.3 Out of Distribution Evaluation
Finally, we evaluate the [Tran et al., 2022] method with different priors and varying numbers of hidden layers on an out-of-distribution (OOD) task. Since the true data distribution in OOD regions is unknown, we rely on optimal transport metrics—1-Wasserstein ($\mathcal{W}_{1}^{1}$) and 2-Wasserstein ($\mathcal{W}_{2}^{2}$). Additionally, we compare all these configurations against our single-hidden-layer method.
@@ E.3 Out of Distribution Evaluation
The results, presented in Fig. 17, 18, 19, and 20, lead to several key observations regarding the performance on OOD tasks:
@@
  • Our method: Due to its grounding in theoretical properties, our approach consistently outperforms all configurations of the [Tran et al., 2022] method.
@@
Figure 16: We compare our method with Tran et al. [2022] on the out of distribution data. For this aim, we provide a wide experimental setting comparing both methods on a few different ranges of data ($x$). We present these ranges as different colour lines in this plot. Firstly, we consider the original test range (R0), then we focus on the data between the given blobs of points (R3). Finally, we consider the out of distribution ranges – ones being closer to the train data (R2, R4, and R2 $\bigcup$ R4) or further (R1, R5, and R1 $\bigcup$ R5). The densities of evaluation points meets the density of the original data on each range despite R0.
Figure 17: Comparison of Tran et al. [2022] method with a Gaussian prior and different number of hidden layers. In this plot, we present $\mathcal{W}_{1}^{1}$ (top row) and $\mathcal{W}_{2}^{2}$ (bottom row) for the regions R3, R4 $\bigcup$ R2, and R1 $\bigcup$ R5 from left to right side. The appropriate regions might be found in Fig. 16.
Figure 18: Comparison of Tran et al. [2022] method with a Gaussian prior and different number of hidden layers. In this plot, we present $\mathcal{W}_{1}^{1}$ (top row) and $\mathcal{W}_{2}^{2}$ (bottom row) for the regions R1, R2, R4, and R5 from left to right side. The appropriate regions might be found in Fig. 16.
Figure 19: Comparison of Tran et al. [2022] method with a Normalizing Flow prior and different number of hidden layers. In this plot, we present $\mathcal{W}_{1}^{1}$ (top row) and $\mathcal{W}_{2}^{2}$ (bottom row) for the regions R3, R4 $\bigcup$ R2, and R1 $\bigcup$ R5 from left to right side. The appropriate regions might be found in Fig. 16.
Figure 20: Comparison of Tran et al. [2022] method with a Normalizing Flow prior and different number of hidden layers. In this plot, we present $\mathcal{W}_{1}^{1}$ (top row) and $\mathcal{W}_{2}^{2}$ (bottom row) for the regions R1, R2, R4, and R5 from left to right side. The appropriate regions might be found in Fig. 16.